Chloe is two years older than her husband Damon. If the product of their ages is thirty more than fifteen times the sum of their ages. How old is chloe?
Answer by Maths68(1474)

Chole’s age = c
Husband’s age = h
Product of their ages = c*h
Sum of their ages = c+h
15 times the sum of their ages = 15 (c+h)
Chole two years older than her husband so if we add 2 years to the husband’s age both will be the same age
c=h+2 …….(1)
the product of their ages is thirty more than fifteen times the sum of their ages
c*h=30+15(c+h)……..(2)
Put the value of c from (1) to (2)
(h+2)*h=30+15[(h+2)+h]
h^2+2h=30+15h+30+15h
h^2+2h=60+30h
h^2+2h-60-30h=0
h^2-28h-60=0
h^2-30h+2h-60=0
h(h-30)-2(h-30)=0
(h+2)(h-30)
h=-2 h=30
h=-2 impossible
so h=30
Husband is 30 years old
And
Chole is 32